Express the system of differential equations in matrix notation x – 4x + y - (cos t)x = 0 y+y - t?x + 3y+e-2x = 0 Which o
X1 X1 Х2 Х2 ОА. Х3 || Х3 X4 X4 X5 X5 х1 81 Х2 Х2 СВ. Х3 Х3 X4 X4
Answer #1

The solution to the above problem is given below. Please review the answer and let me know if you have any doubts.Soi? - 4x+ ty - (cost)0 y! +y-ty+3y +é2t 20 let 04=4 → ble = se =xz] -0 4 = x x = x 713 = y + y = y = H4 |- Hu=yl